OpenStudy (anonymous):

please help! find the exact solutions of the given equation in the interval [0,2pi) sin^2x+sin2x=-cos^2x+1

OpenStudy (anonymous):

add \(\cos^2(x)\) to both sides and then note that \(\sin^2(x)+\cos^2(x)=1\) greatly simplifying the equation

OpenStudy (anonymous):

i take it the original equation is \[\sin^2(x)+\sin(2x)=-\cos^2(x)+1\] right?

OpenStudy (primeralph):

Either way, your first step is right. @satellite73

OpenStudy (anonymous):

yeah thats the equation

OpenStudy (anonymous):

thank you @\(\large' ralph\)

OpenStudy (primeralph):

You remembered.

OpenStudy (primeralph):

@ginger994 Do you know what to do next?

OpenStudy (anonymous):

yeah i know the gist of it, thanks

OpenStudy (anonymous):

you should end up with \[1+\sin(2x)=1\] as a second step

OpenStudy (anonymous):

i got 0, pi/2, pi, 3pi/2 as an answer but i dont think thats right...

OpenStudy (anonymous):

you get \(\sin(2x)=0\) right?

OpenStudy (anonymous):

oh wait, maybe they are ...

OpenStudy (anonymous):

they look good to me

OpenStudy (anonymous):

ok because i was messing around and got 0, pi/3, 2pi/3, pi, 4pi/3 too @satellite73

OpenStudy (anonymous):

no that isn't right, your first answer was right

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

how did you like my rebus?

OpenStudy (primeralph):

@satellite73 Was that picture posted in honor of me?

OpenStudy (anonymous):

of course; it is "primer alf"

OpenStudy (primeralph):

Well, now I know how your childhood went. I'm truly sorry; really I am.

OpenStudy (anonymous):

it had a very productive childhood, it is my adulthood that is being misspent

OpenStudy (primeralph):

@satellite73 You know I was joking right?

OpenStudy (primeralph):

What I meant is that you read really old books as a child. That book is the older version of a children's learning book.

OpenStudy (anonymous):

yes (i hope)